2. the equation N=50+50Y gives the number N of stamps collected when Y is the years that passed since the collection began. Using this equation, determine the number os stamps that will be in the collection after 3 years.
3. The slope of the equation in the above problem is.
4. the equation y=240x+720 gives the cost y (in dollars) when x gadgets are produce. How many gadgets can be produce for $30,000?

Answer by stanbon(75887)

the equation N(Y) = 50+50Y gives the number N of stamps collected when Y is the years that passed since the collection began.
Using this equation, determine the number of stamps that will be in the collection after 3 years.
N(3) = 50 + 50*3 = 5 + 150 = 200 stamps
--------------------------------------------------
3. The slope of the equation in the above problem is.
Ans: 50 because it is the coefficient of Y.
--------------------------------------------------
4. the equation y=240x+720 gives the cost y (in dollars) when x gadgets are produce. How many gadgets can be produced for $30,000?
30000 = 240x + 720
240x = 29280
x = 122 gadgets
=====================
Cheers,
Stan H.